### v1

@@ 5.2 The generalized Heron problem in Hadamard manifolds
In 2012, Mordukhovich et al. [55] generalized the Heron problem. In the examination of the generalized Heron problem, as discussed in reference [55, 56, 2, 3, 5], the conventional notions of points and straight lines are substituted with nonempty, closed and convex sets. Within this context, the objective is to locate a point within a predetermined nonempty, closed and convex set $C\subset\mathbb{R}^{m}$ in such a way that the combined distance to specified nonempty, closed and convex sets $C_{k}\subset\mathbb{R}^{m},k\in\varLambda_{N}$ is minimized. Mathematically, we can write
@@
The constrained problem (59) can be formulated into the following constrained problem:
@@
$\min_{\textbf{x}\in\mathcal{M}^{N+1}}F(\textbf{x})+\iota_{\textbf{D}}(\textbf{x}),$ (62)
[…]
@@
$G(x)=\operatorname{diag}(1/x_{1}^{2},\ldots,1/x_{m}^{2})\text{ and }H(x)=\operatorname{diag}(x_{1},\ldots,x_{m})$ (63)
@@
$\langle u,v\rangle_{R}=\langle u,G(x)v\rangle\text{ for }x\in\mathcal{M}\text{ and }u,v\in\mathcal{T}_{x}\mathcal{M}=\mathbb{R}^{m}.$ (64)
The Riemannian distance $d\colon\mathcal{M}\times\mathcal{M}\to[0,\infty)$ is defined by
@@
The sectional curvature of the Riemannian manifold $\mathcal{M}$ is 0. Thus $M$ is a Hadamard manifold (see [53]). For $x=(x_{1},\ldots,x_{m}),y=(y_{1},\ldots,y_{m})\in\mathcal{M}$, the geodesic joining $x$ to $y$ is defined by
@@
the exponential map $\exp_{\mathbf{x}}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ is defined by
@@
the inverse of the exponential map $\exp_{x}^{-1}\colon\mathcal{M}\to\mathcal{T}_{x}\mathcal{M}$ is defined by
$\exp^{-1}_{x}y=\left(x_{1}\ln\left({y_{1}}/{x_{1}}\right),\ldots,x_{m}\ln\left({y_{m}}/{x_{m}}\right)\right)\text{ for all }y=(y_{1},\ldots,y_{m})\in\mathcal{M}.$ (65)
We denote the closed ball centered at $c\in\mathcal{M}$ with radius $r$ by $B_{r}[c]$, i.e.,
$B_{r}[c]=\{x\in\mathcal{M}:d(x,c)\leq r\}.$
@@
    $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x})=P_{\textbf{D}}(\textbf{x})=\left(b,\ldots,b\right)\in\mathcal{M}^{N}$ where $\textbf{x}=(x_{1},\ldots,x_{N})$, $x_{k}=(x_{k,1},x_{k,2},\ldots,x_{k,m})$, $k\in\varLambda_{N}$ and $b=\left(\left(\prod_{k\in\varLambda_{N}}x_{k,1}\right)^{1/N},\ldots,\left(\prod_{k\in\varLambda_{N}}x_{k,m}\right)^{1/N}\right)$.
@@
    $\operatorname*{Prox}_{\lambda d(x,c)}(x)=\begin{cases}\gamma(x,c;\theta)&\text{ if }d(x,c)>\lambda,\\ c,&\text{ if }d(x,c)\leq\lambda,\end{cases}$ where $\theta=\frac{\lambda}{d(x,c)}$.
@@
    $\operatorname*{Prox}_{\lambda d(x,B_{r}[c])}(x)=\begin{cases}\gamma(x,P_{B_{r}[c]}(x);\theta)&\text{ if }d(x,B_{r}[c])>\lambda,\\ P_{B_{r}[c]}(x),&\text{ if }d(x,B_{r}[c])\leq\lambda,\end{cases}$ where $\theta=\frac{\lambda}{d(x,B_{r}[c])}$.
@@
We compute the solution of the generalized Heron problem using parallel Douglas-Rachford method [4, Algorithm 4], inertial parallel Douglas-Rachford method (58) and $p$-accelerated normal S-parallel Douglas-Rachford method (57) and compare based on their efficiency. Parameters for these algorithms are listed below:
[4, Algorithm 4][Alg(PDRA)] $:$ $\alpha_{n}=0.7;$
Algorithm (58)[Alg(In-M)] $:$ $\alpha_{n}=0.7,\theta_{n}=0.08;$
Algorithm (57)[Alg($p$-Acc)] $:$ $\alpha_{n}=0.7,\beta_{n}=0.5.$
@@
Consider the generalized Heron problem (66), where $m=2,N=2$, the constraint set $C=B_{r}[c]$ is a ball in $\mathcal{M}$ centered at $c=[35~35]^{T}$ with radius $r=0.4$. Consider two cases:
Case 1: Let $a_{1}=[15~70]^{T}$ and $a_{2}=[70~15]^{T}$.
Case 2: Let $a_{1}=[5~50]^{T}$ and $a_{2}=[50~5]^{T}$.
Note that in both cases, $C$ does not contain $c_{1}$ and $c_{2}$. Now we have two observations:
  1. 1.
    In Case 1, the geodesic segment $\gamma(c_{1},c_{2};\cdot)$ intersects the ball $C$. Then any point in the intersection is a solution of problem (66). In this case, the problem (66) has infinitely many solutions.
  2. 2.
    In Case 2, the geodesic segment $\gamma(c_{1},c_{2};\cdot)$ does not intersect the ball $C$. Then, there is a unique point on the boundary of the ball $C$.
@@
We take the initial point $\textbf{x}_{0}=(1,1,1,1)$ for all the algorithms and $\textbf{x}_{1}=(2,2,2,2)$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 40. The performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 2. We have plotted the graph between the number of iterations and $E(n)=d(t_{n},t_{n-1})$ to have a clear visualization of the experiment, where $t_{n}$ is the first $2$ coordinates of $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x}_{n})$; see Figure 4. From Table 2 and Figure 4, we see that our Algorithm Alg($p$-Acc) takes less time and less number of iterations to achieve the desired stopping criteria $E(n)\leq 10^{-10}$.
Table 2: Number of iterations required to have different accuracy for different algorithms with $E(n)<10^{-10}$ as stopping criteria for Example 40
@@
Algorithm CPU(s) iter(n) $E(n)$ CPU(s) iter(n) $E(n)$
@@
Figure 4: Graphical comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $E(n)$ for Example 40.
@@
Case 1: [$(m=2,N=4)$]. We take predetermined set $C$ as ball $B_{r}[c]$ with
$c=(35,35)\text{ and }r=0.4$
and choose constraints sets to be points
$a_{1}=(15,15),a_{2}=(65,65),a_{3}=(10,60)\text{ and }a_{4}=(60,10).$
See Figure 5.
Case 2: [$(m=20,N=10)$]. We take predetermined set $C$ as ball $B_{r}[c]$ with
@@
Figure 5: The generalized Heron problem for four points with ball constraint
We take the initial point $\textbf{x}_{0}=ones(1,m*N)$ for all the algorithms and $\textbf{x}_{1}=2*ones(1,m*N)$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 41. The performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 3 for both cases (Case 1 and Case 2) in Example 41. We have also plotted the graph between the number of iterations and $Er(n)=d(t_{n},t_{n-1})$ to have a clear visualization of the experiment, where $t_{n}$ is the first $2$ coordinates of $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x}_{n})$ in Case 1 and first $20$ coordinates of $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x}_{n})$ in Case 2; see Figure 6. From Table 3 and Figure 6, we see that our Algorithm Alg($p$-Acc) takes less time and less number of iterations to achieve the desired stopping criteria $E_{n}\leq 10^{-10}$ in both cases. From Figure 7, we can see the convergence behaviour of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) to achieving the desired goal as number of iteration increases upto $30$th iteration for Case 2.
Table 3: Number of iterations required to have different accuracy for different algorithms with $Er(n)<10^{-10}$ as stopping criteria for Example 41
@@
Figure 6: Graphical behavior of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) via the graph between the number of iterations and $Er(n)$ for Case 1 and Case 2 in Example 41.
@@
Figure 7: Graphical behavior of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) via the graph between the number of iterations and $Er(n)$ for Case 2 in Example 41.
@@
Case 1: [$(m=2,N=4)$]. We take predetermined set $C$ as ball $B_{r}[c]$ with
$c=(35,35)\text{ and }r=0.4$
and choose constraints sets to be balls $C_{k}=B_{r}[c_{k}]$, $k=1,2,3,4$ with the center
$a_{1}=(15,15),a_{2}=(65,65),a_{3}=(10,60)\text{ and }a_{4}=(60,10),$
respectively and $r=0.4$ See Figure 8 (a).
Case 2: [$(m=3,N=4)$]. We take predetermined set $C$ as ball $B_{r}[c]$ with
$c=(35,35,35)\text{ and }r=0.4$
and choose constraints sets to be balls $C_{k}=B_{r}[c_{k}]$, $k=1,2,3,4$ with the center
$a_{1}=(15,15,15),a_{2}=(65,65,65),a_{3}=(10,60,10)\text{ and }a_{4}=(60,10,60),$
respectively and $r=0.4$. See Figure 8 (b).
@@
We take the initial point $\textbf{x}_{0}=ones(1,m*N)$ for all the algorithms and $\textbf{x}_{1}=2*ones(1,m*N)$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 42. The performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 4 for both cases (Case 1 and Case 2) in Example 42. We have also plotted the graph between the number of iterations and $Er(n)=d(t_{n},t_{n-1})$ to have a clear visualization of the experiment, where $t_{n}$ is the first $2$ coordinates of $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x}_{n})$ in Case 1 and first $20$ coordinates of $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x}_{n})$ in Case 2, see Figure 9. From Table 4 and Figure 9, we see that our Algorithm Alg($p$-Acc) takes less time and less number of iterations to achieve the desired stopping criteria $E_{n}\leq 10^{-10}$ in both cases.
Table 4: Number of iterations required to have different accuracy for different algorithms with $Er(n)<10^{-10}$ as stopping criteria for Example 42
@@
Figure 9: Graphical behavior of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) via the graph between the number of iterations and $Er(n)$ for Case 1 and Case 2 in Example 42.

### v2

@@ 5.2 The generalized Heron problem in Hadamard manifolds
In 2012, Mordukhovich et al. [Mordukhovich2012] generalized the Heron problem. In the examination of the generalized Heron problem, as discussed in reference [Mordukhovich2012, Mordukhovich2012(2), Bot2015, Dixit2022, ShikherOpt2024, ReichFT2023], the conventional notions of points and straight lines are substituted with nonempty, closed and convex sets. Within this context, the objective is to locate a point within a predetermined nonempty, closed and convex set $C\subset\mathbb{R}^{m}$ in such a way that the combined distance to specified nonempty, closed and convex sets $C_{k}\subset\mathbb{R}^{m},k\in\varLambda_{N}$ is minimized. Mathematically, we can write
@@
The constrained problem (59) can be formulated into the following unconstrained problem:
@@
$\min_{\mathbf{x}\in\mathcal{M}^{N+1}}F(\mathbf{x})+\iota_{\mathbf{D}}(\mathbf{x}),$ (62)
where $F(\mathbf{x})=\sum_{k\in\varLambda_{N}}d(x_{k},C_{k})+\iota_{C}(x_{N+1})$ and $\mathbf{D}$$\{\mathbf{x}\in\mathcal{M}^{N+1}\colon x_{1}=\dots=x_{N+1}\}$.
Remark 34.
By (56), the reflection of $F$ defined in (62) is given by
$R_{\lambda F}(\mathbf{x})=\bigl(R_{\lambda d(x_{1},C_{1})}(x_{1}),\ldots,R_{\lambda d(x_{N},C_{N})}(x_{N}),R_{\lambda\iota_{C}}(x_{N+1})\bigr)\quad\text{for all }\mathbf{x}\in\mathcal{M}^{N+1}.$
Since $\mathcal{M}^{N+1}$ has constant sectional curvature, it follows from Proposition 13 and Theorem 14 that the reflection $R_{\lambda F}$ is nonexpansive. Similarly, by Proposition 13, the reflection $R_{\lambda\iota_{\mathbf{D}}}$ is nonexpansive, since $\mathbf{D}$ is a nonempty, closed, and geodesically convex set.
@@
$G(x)=\operatorname{diag}(1/x_{1}^{2},\ldots,1/x_{m}^{2})\text{ and }H(x)=\operatorname{diag}(x_{1},\ldots,x_{m})$
@@
$\langle u,v\rangle_{x}=\langle u,G(x)v\rangle\text{ for }x\in\mathcal{M}\text{ and }u,v\in\mathcal{T}_{x}\mathcal{M}=\mathbb{R}^{m}.$ (63)
The Riemannian distance $d\colon\mathcal{M}\times\mathcal{M}\to[0,\infty)$ is given by
@@
$\mathcal{M}$ is a Hadamard manifold with zero sectional curvature (see [DaCruz]). For $x=(x_{1},\ldots,x_{m}),y=(y_{1},\ldots,y_{m})\in\mathcal{M}$, the geodesic joining $x$ to $y$ is given by
@@
the exponential map $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ is given by
@@
the inverse of the exponential map $\exp_{x}^{-1}\colon\mathcal{M}\to\mathcal{T}_{x}\mathcal{M}$ is given by
$\exp^{-1}_{x}y=\left(x_{1}\ln\left({y_{1}}/{x_{1}}\right),\ldots,x_{m}\ln\left({y_{m}}/{x_{m}}\right)\right)\text{ for all }y=(y_{1},\ldots,y_{m})\in\mathcal{M}.$
We denote the closed ball centered at $c\in\mathcal{M}$ with radius $r$ by $B_{r}[c]$, i.e., $B_{r}[c]=\{x\in\mathcal{M}:d(x,c)\leq r\}.$
@@
    $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x})=P_{\mathbf{D}}(\mathbf{x})=\left(b,\ldots,b\right)\in\mathcal{M}^{N}$ for all $\mathbf{x}=(x_{1},\ldots,x_{N})\in\mathcal{M}^{N}$, where $x_{k}=(x_{k,1},x_{k,2},\ldots,x_{k,m})$, $k\in\varLambda_{N}$ and $b=\left(\left(\prod_{k\in\varLambda_{N}}x_{k,1}\right)^{\frac{1}{N}},\ldots,\left(\prod_{k\in\varLambda_{N}}x_{k,m}\right)^{\frac{1}{N}}\right)$.
@@
    $\operatorname*{Prox}_{\lambda d(x,c)}(x)=\begin{cases}\gamma(x,c;\theta),&\text{ if }d(x,c)>\lambda,\\ c,&\text{ if }d(x,c)\leq\lambda,\end{cases}$ where $\theta=\frac{\lambda}{d(x,c)}$.
@@
    $\operatorname*{Prox}_{\lambda d(x,B_{r}[c])}(x)=\begin{cases}\gamma(x,P_{B_{r}[c]}(x);\theta),&\text{ if }d(x,B_{r}[c])>\lambda,\\ P_{B_{r}[c]}(x),&\text{ if }d(x,B_{r}[c])\leq\lambda,\end{cases}$ where $\theta=\frac{\lambda}{d(x,B_{r}[c])}$.
@@
We compute the solution of the generalized Heron problem using parallel Douglas-Rachford method [Bergmann2016, Algorithm 4], inertial parallel Douglas-Rachford method (58) and $p$-accelerated normal S-parallel Douglas-Rachford method (57) and compare based on their efficiency.
By using Remark 36(a), we may assume that $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x}_{n})=(t_{n},\ldots,t_{n})\in\mathcal{M}^{N+1}$. We then define the error term by $Er(n)=d(t_{n},t_{n-1})$. In all experiments, the algorithm is terminated when the stopping criterion $Er(n)\leq 10^{-10}$ is satisfied.
We take $\lambda=1$ and iteration parameters for these algorithms are listed below:
[Bergmann2016, Algorithm 4][Alg(PDRA)] $:$ $\alpha_{n}=0.4;$
Algorithm (58)[Alg(In-M)] $:$ $\alpha_{n}=0.4,\theta_{n}=0.08;$
Algorithm (57)[Alg($p$-Acc)] $:$ $\alpha_{n}=0.4.$
Note that we use the same parameter $\alpha_{n}$ for all the algorithms. However, the inertial parameter $\theta_{n}$ is chosen according to condition (C3), that is, $\epsilon=(1-\alpha_{n})/\alpha_{n}$, $\epsilon_{1}=1+\epsilon+\max\{1,\epsilon\}$, and $\theta_{n}\approx 0.25(\epsilon/\epsilon_{1})$. We chose $p=2$, which is the smallest value that yielded the best performance in our experiments. The parameters were chosen by testing several reasonable values, and those giving the best convergence are reported.
@@
Consider the generalized Heron problem (64), where $m=2,N=2$, the constraint set $C=B_{r}[c]$ is a ball in $\mathcal{M}$ centered at $c=(35,35)$ with radius $r=0.4$. Consider two cases:
Case 1: Let $a_{1}=(15,70)$ and $a_{2}=(70,15)$.
Case 2: Let $a_{1}=(5,50)$ and $a_{2}=(50,5)$.
Note that in both cases, $C$ does not contain $a_{1}$ and $a_{2}$. Now we have two observations:
  1. (i)
    In Case 1, the geodesic segment $\gamma(a_{1},a_{2};\cdot)$ intersects the ball $C$; see Figure 3(a). Then any point in the intersection is a solution of problem (64). In this case, problem (64) has infinitely many solutions.
  2. (ii)
    In Case 2, the geodesic segment $\gamma(a_{1},a_{2};\cdot)$ does not intersect the ball $C$; see Figure 3(b). Then, there is a unique point on the boundary of the ball $C$.
@@
We take the initial point $\mathbf{x}_{0}=ones(1,6)$ for all the algorithms and $\mathbf{x}_{1}=2*ones(1,6)$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 37. The numerical performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 3. We plot the relationship between the number of iterations and $Er(n)$; see Figure 4. From Table 3 and Figure 4, we observe that Alg($p$-Acc) requires less computational time and fewer iterations to achieve the desired stopping criterion in both cases.
Table 3: Numerical performance in terms of CPU time and number of iterations for Alg(PDRA), Alg(In-M), and Alg($p$-Acc) for Example 37
@@
Algorithm CPU(s) iter(n) $Er(n)$ CPU(s) iter(n) $Er(n)$
@@
Figure 4: Comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $Er(n)$ for Example 37.
@@
Case 1: $m=2$, $N=4$.
We take predetermined set $C$ as ball $B_{r}[c]$ with $c=(35,35)$ and $r=0.4$ and choose constraints sets to be points $a_{1}=(15,15)$, $a_{2}=(65,65)$, $a_{3}=(10,60)$ and $a_{4}=(60,10)$; see Figure 5.
Case 2: $m=20$, $N=10$.
We take predetermined set $C$ as ball $B_{r}[c]$ with
@@
Figure 5: The generalized Heron problem for four points with a ball constraint
We take the initial point $\mathbf{x}_{0}=ones(1,m*(N+1))$ for all the algorithms and $\mathbf{x}_{1}=2*ones(1,m*(N+1))$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 38. The numerical performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 4 for both cases (Case 1 and Case 2). We plot the relationship between the number of iterations and $Er(n)$ to have a clear visualization of the experiment; see Figure 6. From Table 4 and Figure 6, we see that our Algorithm Alg($p$-Acc) requires less computational time and less number of iterations to achieve the desired stopping criterion in both cases. From Figure 7, we can see the convergence behaviour of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) to achieving the desired goal as number of iteration increases upto $30$th iteration for Case 2.
Table 4: Numerical performance in terms of CPU time and number of iterations for Alg(PDRA), Alg(In-M), and Alg($p$-Acc) for Example 38
@@
Figure 6: Comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $Er(n)$ for Example 38.
@@
Figure 7: Graphical progression of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) upto 30 iterations for Case 2 in Example 38.
@@
Case 1: $m=2$, $N=4$.
We take predetermined set $C$ as ball $B_{r}[c]$ with $c=(35,35)$ and $r=0.4$ and choose constraints sets to be balls $C_{k}=B_{r}[a_{k}]$, $k=1,2,3,4$ with the center $a_{1}=(15,15)$, $a_{2}=(65,65)$, $a_{3}=(10,60)$ and $a_{4}=(60,10)$ and $r=0.4$; see Figure 8(a).
Case 2: $m=3$, $N=4$.
We take set $C$ as ball $B_{r}[c]$ with $c=(35,35,35)$ and $r=0.4$ and choose constraints sets to be balls $C_{k}=B_{r}[a_{k}]$, $k=1,2,3,4$ with the center $a_{1}=(15,15,15)$, $a_{2}=(65,65,65)$, $a_{3}=(10,60,10)$ and $a_{4}=(60,10,60)$ and $r=0.4$; see Figure 8(b).
@@
We take the initial point $\mathbf{x}_{0}=ones(1,m*(N+1))$ for all the algorithms and $\mathbf{x}_{1}=2*ones(1,m*(N+1))$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 39. The numerical performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 5 for both cases (Case 1 and Case 2). We plot the relationship between the number of iterations and $Er(n)$ to have a clear visualization of the experiment; see Figure 9. From Table 5 and Figure 9, we see that our Algorithm Alg($p$-Acc) takes less runtime and less number of iterations to achieve the desired stopping criterion in both cases.
Table 5: Numerical performance in terms of CPU time and number of iterations for Alg(PDRA), Alg(In-M), and Alg($p$-Acc) for Example 39
@@
Figure 9: Comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $Er(n)$ for Example 39.